# MCDubber: Multimodal Context-Aware Expressive Video Dubbing

### v1

@@ Abstract
Current AVD methods generally focus two main issues. 1) Duration control and audio-visual synchronization. Neural Dubber [1] and DSU-AVO [7] employ a text-video aligner by cross-modal attention to learn the single sentence alignment between lip motion and text, aiming to address these challenges. 2) Prosody expressiveness in video dubbing. VDTTS [8], HPMDubbing [9], and 3D-VD [2] improve the prosody expressiveness of dubbing by integrating facial visual information of current (target) sentence. These methods are sometimes limited by their insufficient consideration of the multi-modal context of the current sentence being dubbed.
@@ 5.2 Results of Video Dubbing
To evaluate the prosody expressiveness of MCDubber, we develope three baseline models as following. (1) FastSpeech2 [6]: a neural TTS model; (2) DSU-AVO [7], an AVD model with a learning objective focused on discrete speech unit prediction; (3) HPMDubbing [9], an AVD model incorporating hierarchical prosody modeling considering facial expressions. For a fair comparison, these baseline models were trained on the Original Chem dataset. Note that all of the above baseline models are designed for modeling a single sentence.
The results, shown in Table 3, indicate that our method achieves the best performance on prosody-related metrics. Specifically, our method attains a GPE of 40.08 and an FFE of 31.74, surpassing previous methods. This demonstrates that our model generates more prosody expressive speech. Additionally, our method also achieved the best performance in prosody-related metrics, specifically MOS-S and MOS-C, as determined by human subjective evaluations. MOS-S shows our synthesized speech is closer to ground-truth speech in terms of prosody. MOS-C indicates that after integrating the dubbed video with the original context, the prosody of the generated dubbing aligns well with the multimodal context, outperforming other models. In summary, both objective and subjective metrics demonstrate that our model improves dubbing prosody expressiveness and enhances the prosody consistency of the dubbed video with the multimodal context compared to other methods.
@@
| GT | N/A | N/A | N/A | 8.722 | 6.565 | 4.23 ± 0.08 | 4.39 ± 0.07 | 4.41 ± 0.07 | 4.44 ± 0.07 | 4.46 ± 0.07 |
| Mel Resynthesis | 02.66 | 11.58 | 0.952 | 8.482 | 6.551 | 4.14 ± 0.05 | 4.20 ± 0.07 | 4.21 ± 0.05 | 4.27 ± 0.07 | 4.24 ± 0.06 |
| FastSpeech2 | 50.26 | 41.85 | 0.211 | 3.282 | 9.414 | 3.50 ± 0.12 | 3.36 ± 0.08 | 3.53 ± 0.06 | 3.39 ± 0.11 | 3.44 ± 0.12 |
| DSU-AVO | 48.02 | 31.75 | 0.536 | 8.198 | 7.119 | 3.77 ± 0.06 | 3.74 ± 0.06 | 3.55 ± 0.06 | 3.72 ± 0.10 | 3.90 ± 0.10 |
| HPMDubbing | 42.73 | 38.79 | 0.388 | 7.929 | 7.053 | 3.69 ± 0.05 | 3.81 ± 0.05 | 3.84 ± 0.07 | 3.82 ± 0.08 | 3.91 ±0.07 |
| Ours | 40.08 | 31.74 | 0.508 | 7.566 | 6.941 | 3.83 ± 0.05 | 3.97 ± 0.06 | 4.02 ± 0.06 | 4.03 ± 0.08 | 3.96 ± 0.07 |

### v3

@@ Abstract
Current AVD methods generally focus on two main issues. 1) Duration control and audio-visual synchronization. Neural Dubber [1] and DSU-AVO [7] employ a text-video aligner by cross-modal attention to learn the single sentence alignment between lip motion and text, aiming to address these challenges. 2) Prosody expressiveness in video dubbing. VDTTS [8], HPMDubbing [9], and 3D-VD [2] improve the prosody expressiveness of dubbing by integrating facial visual information of current (target) sentence. These methods are sometimes limited by their insufficient consideration of the multi-modal context of the current sentence being dubbed.
@@ 5.2 Results of Video Dubbing
To evaluate the prosody expressiveness of MCDubber, we develop three baseline models as follows. (1) FastSpeech2 [6]: a neural TTS model; (2) DSU-AVO [7], an AVD model with a learning objective focused on discrete speech unit prediction; (3) HPMDubbing [9], an AVD model incorporating hierarchical prosody modeling considering facial expressions. For a fair comparison, these baseline models were trained on the Original Chem dataset. Note that all of the above baseline models are designed for modeling a single sentence. The results, shown in Table 3, indicate that our method achieves the best performance on prosody-related metrics. Specifically, our method attains a GPE of 40.26 and an FFE of 30.94, surpassing previous methods. This demonstrates that our model generates more prosody expressive speech. Additionally, our method also achieved the best performance in prosody-related metrics, specifically MOS-S and MOS-C, as determined by human subjective evaluations. MOS-S shows our synthesized speech is closer to ground-truth speech in terms of prosody. MOS-C indicates that after integrating the dubbed video with the original context, the prosody of the generated dubbing aligns well with the multimodal context, outperforming other models. In summary, both objective and subjective metrics demonstrate that our model improves dubbing prosody expressiveness and enhances the prosody consistency of the dubbed video with the multimodal context compared to other methods.
@@
| GT | N/A | N/A | N/A | 8.722 | 6.564 | 4.23 ± 0.08 | 4.39 ± 0.07 | 4.41 ± 0.07 | 4.44 ± 0.07 | 4.46 ± 0.07 |
| Mel Resynthesis | 02.77 | 09.70 | 0.952 | 8.468 | 6.538 | 4.14 ± 0.05 | 4.20 ± 0.07 | 4.21 ± 0.05 | 4.27 ± 0.07 | 4.24 ± 0.06 |
| FastSpeech2 | 50.26 | 41.24 | 0.211 | 3.438 | 9.399 | 3.50 ± 0.12 | 3.36 ± 0.08 | 3.53 ± 0.06 | 3.39 ± 0.11 | 3.44 ± 0.12 |
| DSU-AVO | 48.02 | 32.31 | 0.536 | 8.198 | 7.120 | 3.77 ± 0.06 | 3.74 ± 0.06 | 3.55 ± 0.06 | 3.72 ± 0.10 | 3.90 ± 0.10 |
| HPMDubbing | 42.78 | 36.56 | 0.383 | 7.934 | 7.057 | 3.69 ± 0.05 | 3.81 ± 0.05 | 3.84 ± 0.07 | 3.82 ± 0.08 | 3.91 ±0.07 |
| Ours | 40.26 | 30.94 | 0.508 | 7.571 | 6.943 | 3.83 ± 0.05 | 3.97 ± 0.06 | 4.02 ± 0.06 | 4.03 ± 0.08 | 3.96 ± 0.07 |